Need help with the 3 parts of this problem.

Superior Company provided the following data for the year ended December 31 (al raw materials are used in production as direct materials s 218,000 267,000 Selling expenses Direct labor Manufacturing overhead applied to vork in process 365,000 Actual manufacturing overhead cost Inventory balances at the beginning and end of the year were as follows Beginning of Year End of Year $ 54,000 39,000 7 32,000 Work in process Pinished goods The total manufacturing costs for the year were $690,000; the cost of goods available for sale totaled $740,000; the unadjusted cost of goods sold totaled $667,000; and the net operating income was $31,000. The companys underapplied or overapplied overhead is closed to Cost of Goods Sold. Required: Prepare schedules of cost of goods manufactured and cost of goods sold and an income statement. (Hint: Prepare the income statement and schedule of cost of goods sold first followed by the schedule of cost of goods manufactured Complete this question by entering your answers in the tabs below. Income Statement COGS Scheduke COGM Schedule Prepare an income statement for the year
Answer #1
A B C D
Superior Company
Income Statement
4 Sales revenue 1056000
5 Cost of goods sold 656000
6 Gross profit 400000
7 Selling and administrative expenses:
8 Selling expenses 218000
9 Administrative expenses 151000
10
11 369000
12 Net operating income 31000
Superior Company
Schedule of Cost of Goods Sold
16 Cost of goods manufactured 704000
17 Add: Beginning finished goods inventory 36000
18 Cost of goods available for sale 740000
19 Less: Ending finished goods inventory 73000
20 Unadjusted cost of goods sold 667000
21 Less: Overapplied manufacturing overhead (365000 - 354000) 11000
22 Adjusted cost of goods sold 656000
Superior Company
Schedule of Cost of Goods Manufactured
Direct materials:
27 Beginning raw materials inventory 54000
28 Add: Purchases of raw materials 267000
29 Total raw materials available 321000
30 Less: Ending raw materials inventory 39000
31 Raw materials used in production 282000
32 Direct labor 43000
33 Manufacturing overheads 365000
34 Total manufacturing costs for the year 690000
35 Add: Beginning work in process inventory 46000
36 Total work in process 736000
37 Less: Ending work in process inventory 32000
38 Cost of goods manufactured 704000

Working:

A B C D
Superior Company
Income Statement
4 Sales revenue =D5+D6
5 Cost of goods sold 656000
6 Gross profit =D11+D12
7 Selling and administrative expenses:
8 Selling expenses 218000
9 Administrative expenses 151000
10
11 =SUM(C8:C10)
12 Net operating income 31000
Superior Company
Schedule of Cost of Goods Sold
16 Cost of goods manufactured =C18-C17
17 Add: Beginning finished goods inventory 36000
18 Cost of goods available for sale 740000
19 Less: Ending finished goods inventory =C18-C20
20 Unadjusted cost of goods sold 667000
21 Less: Overapplied manufacturing overhead (365000 - 354000) =365000-354000
22 Adjusted cost of goods sold =C20-C21
Superior Company
Schedule of Cost of Goods Manufactured
Direct materials:
27 Beginning raw materials inventory 54000
28 Add: Purchases of raw materials 267000
29 Total raw materials available =C27+C28
30 Less: Ending raw materials inventory 39000
31 Raw materials used in production =C29-C30
32 Direct labor =D34-D31-D33
33 Manufacturing overheads 365000
34 Total manufacturing costs for the year 690000
35 Add: Beginning work in process inventory =D36-D34
36 Total work in process =D38+D37
37 Less: Ending work in process inventory 32000
38 Cost of goods manufactured 704000